Your research team analysis the light of a mysterious object in space. By using a spectrometer,
you can observe the following spectrum of the object. The Hα line peak is clearly visible:
(a) Mark the first four spectral lines of hydrogen (Hα, Hβ, Hγ, Hδ) in the spectrum.
(b) Determine the radial velocity and the direction of the object’s movement.
(c) Calculate the distance to the observed object.
(d) What possible type of object is your team observing?
1
Expert's answer
2020-06-22T11:19:51-0400

(a) See the spectrum below:



(b) Find the radial velocity using the table value for H-alpha wavelength:


"v=cz=(3\cdot10^5)\bigg(\frac{800}{656.28}-1\bigg)=65700\text{ km/s}."

(c) Find the distance applying Hubble's law:


"D=\frac{v}{H_0}=\frac{65700\text{ km/s}}{67.4\text{ km/s/MPc}}=975\text{ MPc}."

(d) Most likely this is a cluster galaxy.